OpenStudy (rockstar0765):

The probability that a given urine sample analyzed in a pathology lab contains protein is 0.73 (event A). The probability that the urine sample contains the protein albumin is 0.58 (event B). The probability that a sample contains protein, given that albumin is detected, is 1. Which statement is true?

OpenStudy (rockstar0765):

Events A and B are dependent because P(A|B) = P(A). Events A and B are dependent because P(A| B) P(B). Events A and B are dependent because P(A|B) = P(A) + P(B). Events A and B are dependent because P(A| B) P(A). Events A and B are dependent because P(A|B) = P(A).

OpenStudy (error1603):

g